# QOG:Question and Options Generation based on Language Model

### v1

@@ Abstract
§ 1 Introduction
Question-Options (QO) is derived from Question-Answering (QA), extending question-answer pairs by adding three text-related distractors. Compared to QA, QO provides more information by combining the correct answer with several specially designed incorrect answers. This information enables the model to incorporate more knowledge and improve its ability to identify potential errors. When used for fine-tuning large language modelHu et al. (2021); Liu et al. (2021, 2023), the goal is to identify the correct answer from four candidate answers, which improves the model’s understanding of the question and its ability to distinguish the differences between different options, making the model understand the text better. When used for model evaluation, whether the model can select the correct answer becomes a key indicator of its text learning and processing capabilities Hendrycks et al. (2020); Huang et al. (2024); Cobbe et al. (2021).
Question and Options Generation (QOG) refers to the task of generating a set of question-options pairs given an input context (e.g. a paragraph). QOG can be used to develop unsupervised question answering and as a data augmentation tool Alberti et al. (2019) to enhance the text understanding capabilities of large language models. QOG can also be used as an educational aid, to improve information retrieval modelsNogueira et al. (2019); Pyatkin et al. (2021), and as a tool for model explanation.
@@
QOG originates from Question Generation (QG)Duan et al. (2017); Du et al. (2017); Ali et al. (2010), which involves generating a question for an answer given an input context. Compared to QG and QAG, QOG is a more complex task because the correct answer and distractors need to be generated rather than assumed as part of the input. Currently, it is unknown which models are effective for the QOG task because no comprehensive comparative study has been established to date.
In this paper, we consider the QOG as a task of generating questions and options given a context, and compare three simple QOG methods based on fine-tuning encoder-decoder language models such as T5Raffel et al. (2020). The three methods are in accordance with Ushio et al. (2023): (1) Pipeline QOG, which decomposes the task into three parts: answer extraction, question generation and distractor generation, and learns a separate model for each subtask; (2) Multi-task QOG, which uses a shared single model to train the three subtasks simultaneously instead of separate models; and (3) End-to-end QOG, which directly generates questions-options using end-to-end sequence-to-sequence learning. Finally, we introduce GPT-4 as a referee based on traditional evaluation methods and compare the three methods in multiple domains to objectively evaluate the cross-domain generalization ability of the models.
§ 2 Related Work
@@
§ 2.1 Question Generation
Question generation (QG) aims to generate questions given documents and answers, which has been applied in various scenarios such as data augmentationAlberti et al. (2019) and document retrievalNogueira et al. (2019). Puri et al. (2020) fine-tuned QG’s autoregressive LM to generate related questions. These works focus on generating large-scale data at once without paying much attention to data quality. Xiong and Wu (2020) followed the idea of adversarial generation and used question-answering agents and QG models to train adversarial models. This work used semantic similarity as an evaluation indicator to improve the quality of generated data, but did not focus on the generalization ability of the model in other fields. To this end, our work will focus on improving data quality while testing the generalization ability of the model with data sets from different fields.
§ 2.2 Question-Answer Generation
Question-Answer Generation (QAG) generates question-answer pairs based on a given document. QAG can be used to develop QA models without human supervision and as an auxiliary tool for data augmentation of large language models. Bartolo et al. (2021) used QAG model to generate adversarial examples for QA, Lewis et al. (2021) improved extractive QA by generating millions of question-answer pairs through QAG, In both examples, the fine-tuned model is BART. In order to improve the recall rate of QA, Fang et al. (2020)regarded answer extraction as a classification problem and used RoBERTaLiu et al. (2019) to implement it. Han et al. (2023) generated QA iteratively and improved the data quality of each iteration by modifying the seed set. Ushio et al. (2023) made a comparison of several QAG methods and established a complete system. These works can generate high-quality QA pairs, however they are limited to machine metrics and do not evaluate QA quality according to human standards.
§ 2.3 QG & QAG Evaluation
@@
§ 3 Question & Options Generation
@@
$\tilde{a}$ $=\operatorname*{arg\,max}_{a}P_{{\rm ae}}(a|c,s)$ (1)
$\tilde{q}$ $=\operatorname*{arg\,max}_{q}P_{{\rm qg}}(q|c,s,a)$ (2)
$\tilde{d}$ $=\operatorname*{arg\,max}_{d}P_{{\rm dg}}(d|c,s,a)$ (3)
@@
where $c_{i}$ is the $i$-th token in context $c$, and $|n|$ represents the number of tokens in the text. When input to the QG model, the answer $a$ is considered and the position of the answer $a$ in the context is marked with <hl>. This is the work of Ushio et al. (2022). Its form is:
@@
$\mathcal{T}(\mathcal{Q}_{c})=``\texttt{question:}q_{1},\texttt{options:}o_{1}\texttt{|}$
[…]
@@
$\tilde{y}=\operatorname*{arg\,max}_{y}P_{{\rm qog}}(y|c)$ (6)
§ 4 Evaluation
§ 4.1 Experimental Setting
@@
§ 4.3 Assessing the quality of QO pairs using GPT-4
@@
§ 4.4 QOG Model Comparison
(a) Inference time
(b) Memory usage
Figure 3: Inference time and memory usage for different QOG models
@@
§ 5 Conclusion
@@
Contributions
This paper establishes the English dataset SQuAD-QO for training QOG, establishes a baseline of QOG, and also provides a feasible evaluation idea.

### v2

@@ Abstract
keywords:
Question-Options Generation (QOG); Information retrieval; sequence-to-sequence language models
editors: Vu Nguyen and Hsuan-Tien Lin
1 Introduction
Question-Options (QO) is derived from Question-Answering (QA), extending question-answer pairs by adding three text-related distractors. Compared to QA, QO provides more information by combining the correct answer with several specially designed incorrect answers. This information enables the model to incorporate more knowledge and improve its ability to identify potential errors. When used for fine-tuning large language model(Hu et al., 2021; Liu et al., 2021, 2023), the goal is to identify the correct answer from four candidate answers, which improves the model’s understanding of the question and its ability to distinguish the differences between different options, making the model understand the text better. When used for model evaluation, whether the model can select the correct answer becomes a key indicator of its text learning and processing capabilities (Hendrycks et al., 2020; Huang et al., 2024; Cobbe et al., 2021).
Question and Options Generation (QOG) refers to the task of generating a set of question-options pairs given an input context (e.g. a paragraph). QOG can be used to develop unsupervised question answering and as a data augmentation tool (Alberti et al., 2019; Yu et al., 2018; Riabi et al., 2020; Singh et al., 2019; Longpre et al., 2019) to enhance the text understanding capabilities of large language models. QOG can also be used as an educational aid(Agarwal et al., 2019; Cai et al., 2023; Wang et al., 2023), to improve information retrieval models(Nogueira et al., 2019; Pyatkin et al., 2021), and as a tool for model explanation.
@@ 1 Introduction
QOG originates from Question Generation (QG)(Duan et al., 2017; Du et al., 2017; Ali et al., 2010), which involves generating a question for an answer given an input context. Compared to QG and QAG, QOG is a more complex task because the correct answer and distractors need to be generated rather than assumed as part of the input. Currently, it is unknown which models are effective for the QOG task because no comprehensive comparative study has been established to date.
In this paper, we consider the QOG as a task of generating questions and options given a context, and compare three simple QOG methods based on fine-tuning encoder-decoder language models such as T5(Raffel et al., 2020). The three methods are common methods for fine-tuning language models now: (1) Pipeline QOG, which decomposes the task into three parts: answer extraction, question generation and distractor generation, and learns a separate model for each subtask; (2) Multi-task QOG, which uses a shared single model to train the three subtasks simultaneously instead of separate models; and (3) End-to-end QOG, which directly generates questions-options using end-to-end sequence-to-sequence learning. Finally, we introduce GPT-4 as a referee based on traditional evaluation methods and compare the three methods in multiple domains to objectively evaluate the cross-domain generalization ability of the models.
2 Related Work
@@
2.1 Question Generation
Question generation (QG) aims to generate questions given documents and answers, which has been applied in various scenarios such as data augmentation(Alberti et al., 2019; Pellicer et al., 2023) and document retrieval(Nogueira et al., 2019; Hofstätter et al., 2020; Ram et al., 2023). Puri et al. (2020) fine-tuned QG’s autoregressive LM to generate related questions. These works focus on generating large-scale data at once without paying much attention to data quality. Xiong and Wu (2020) followed the idea of adversarial generation and used question-answering agents and QG models to train adversarial models. This work used semantic similarity as an evaluation indicator to improve the quality of generated data, but did not focus on the generalization ability of the model in other fields. To this end, our work will focus on improving data quality while testing the generalization ability of the model with data sets from different fields.
2.2 Question-Answer Generation
Question-Answer Generation (QAG) generates question-answer pairs based on a given document. QAG can be used to develop QA models without human supervision and as an auxiliary tool for data augmentation of large language models. Bartolo et al. (2021) used QAG model to generate adversarial examples for QA, Lewis et al. (2021) improved extractive QA by generating millions of question-answer pairs through QAG, In both examples, the fine-tuned model is BART. In order to improve the recall rate of QA, Fang et al. (2020) regarded answer extraction as a classification problem and used RoBERTa(Liu et al., 2019) to implement it. Han et al. (2023) generated QA iteratively and improved the data quality of each iteration by modifying the seed set. Ushio et al. (2023) made a comparison of several QAG methods and established a complete system. These works can generate high-quality QA pairs, however they are limited to machine metrics and do not evaluate QA quality according to human standards.
2.3 QG & QAG Evaluation
@@
3 Question & Options Generation
@@ 3.1 Pipeline QOG
$\tilde{a}$ $=argmax_{a}P_{{\rm ae}}(a|c,s)$ (1)
$\tilde{q}$ $=argmax_{q}P_{{\rm qg}}(q|c,s,a)$ (2)
$\tilde{d}$ $=argmax_{d}P_{{\rm dg}}(d|c,s,a)$ (3)
@@
where $c_{i}$ is the $i$-th token in context $c$, and $|n|$ represents the number of tokens in the text. When input to the QG model, the answer $a$ is considered and the position of the answer $a$ in the context is marked with <hl>. The form is:
@@ 3.3 End2end QOG
$\mathcal{T}(\mathcal{Q}_{c})=``\texttt{question:}q_{1},\texttt{options:}o_{1}\texttt{|}\texttt{question:}q_{2},\texttt{options:}o_{2}\texttt{|}\dots\textrm{'}\textrm{'}$ ’ ’ (5)
@@
$\tilde{y}=argmax_{y}P_{{\rm qog}}(y|c)$ (6)
4 Evaluation
4.1 Experimental Setting
@@ 4.2 Results
4.3 Assessing the quality of QO pairs using GPT-4
@@
4.4 QOG Model Comparison
Figure 3: Inference time
Figure 4: Memory usage
5 Conclusion
@@
Appendix A Hyper Parameters
In fine-tuning each QOG model, we searched for the optimal hyperparameters, and Table 5 shows the best hyperparameters. The maximum input length is fixed as 512, and the maximum output length is 256.
Approach Model Epoch LR Batch
Pipeline(AE) T5SMALL 6 0.0001 256
Pipeline(QG) T5SMALL 9 0.0001 64
Pipeline(DG) T5SMALL 13 0.00005 64
Multitask T5SMALL 7 0.0001 128
End2end T5SMALL 16 0.0001 128
Pipeline(AE) T5BASE 8 0.0001 64
Pipeline(QG) T5BASE 5 0.0001 32
Pipeline(DG) T5BASE 13 0.00005 32
Multitask T5BASE 7 0.0001 64
End2end T5BASE 17 0.0001 64
Pipeline(AE) T5LARGE 9 0.0001 32
Pipeline(QG) T5LARGE 6 0.00005 32
Pipeline(DG) T5LARGE 8 0.00001 32
Multitask T5LARGE 5 0.0001 32
End2end T5LARGE 12 0.0001 32
Table 5: Optimal hyperparameters for each QOG model.
Appendix B Additional Results of LLM QO evaluation
Table 6 shows the QO evaluation results of some large language models with small parameters. Qwen2-7B has the highest score among all models. Phi-3, as a smaller parameter model, also performs well in the QO evaluation.
Model Aver Amazon Wiki NYT Reddit Fin
Qwen2-7B 57.98 47.37 58.72 55.29 52.51 76.03
Qwen1.5-MoE-A2.7B 48.73 34.16 55.29 43.56 35.91 74.72
Deepseek-7B 54.32 51.58 37.78 61.20 62.09 58.95
GLM-4-9B 42.00 51.56 43.06 8.61 36.24 70.55
Baichuan2-7B 43.64 48.82 33.07 39.01 18.35 78.97
Mistral-7B 38.68 42.86 34.23 37.32 44.45 34.54
Gemma-7B 43.38 48.59 46.67 57.45 10.93 53.26
Phi-3-3.8B 52.98 48.25 52.01 70.48 51.07 43.09
Table 6: QO evaluation results($F_{1}$) of LLM on the test set.
Appendix C Additional Results of GPT-4 evaluation
We also use the GPT evaluation method to evaluate the QO generation quality of each llm. The result is shown in Table 7.
Model GPT-4
Qwen2-7B 98
Qwen1.5-MoE-A2.7B 88
Deepseek-7B 95
GLM-4-9B 81
Baichuan2-7B 85
Mistral-7B 83
Gemma-7B 93
Phi-3-3.8B 87
Table 7: Mean scores for each LLM under the GPT-4 judgement